OpenStudy (anonymous):

What is the distance between the points (–3, 5) and (2, –6)? Round your answer to the nearest tenth. A. 1.4 units B. 5.1 units C. 12.1 units D. 16.0 units

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

i do not understand that

hartnn (hartnn):

then lets take each term separately \(x_1-x_2 = \text{difference between x co-ordinates }=2-(-3)=...? \)

OpenStudy (anonymous):

5?

hartnn (hartnn):

yes! 5 is correct, and in the formula since we square it we get 25. \(y_1-y_2 = \text{difference between y co-ordinates }=5-(-6)=...? \)

OpenStudy (anonymous):

B?

hartnn (hartnn):

5-(-6) would be a number actually

hartnn (hartnn):

and no, its not option B